OpenStudy (anonymous):

HELP HELP PLEASE. find the tangents of the curve 2x62 - 3y^2 + 12x - 12y + 22 = 0 from (7, 6). the answers are 2x - 3y + 4 and 22x - 27y +8 = 0

OpenStudy (anonymous):

what i did was i first get the dy/dx of the equation and it was -12-4x / -12-6y..

OpenStudy (anonymous):

then equate it to the slope form. or the m= y-y1/ x-x1..

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you can simplify this dy/dx = -12-4x / -12-6y \[ \frac{dy}{dx}= \frac{4(x+3)}{6(y+2)} = \frac{2(x+3)}{3(y+2)} \] now do the usual, find the slope between the point (x,y) and (7,6) \[ \frac{2(x+3)}{3(y+2)} = \frac{y-6}{x-7}\]

OpenStudy (anonymous):

typed it wrong. it should be 2x^2.. not 2x62.. oh my @phi thank you so much!!!!!

OpenStudy (phi):

yes, I noticed (because I make the typo often enough)

OpenStudy (anonymous):

got stuck up after cross multiplying..

OpenStudy (anonymous):

2x^2 - 8x - 42= 3y^2 - 12y -36..

OpenStudy (phi):

but you do have some nasty algebra to do... \[ 2(x+3)(x-7) = 3(y+2)(y-6) \\ 2x^2 -8x -42 = 3y^2 -12y -36 \\ (2x^2 - 3y^2) -8x -42 +12y+36=0 \]

OpenStudy (phi):

now we use this trick \[ 2x^2 - 3y^2 + 12x - 12y + 22 = 0 \\2x^2 - 3y^2= -12x+12y-22 \]

OpenStudy (phi):

we can get rid of the 2x^2 - 3y^2, and solve for x

OpenStudy (anonymous):

what do you mean??

OpenStudy (anonymous):

get rid of the 2 terms? why?

OpenStudy (phi):

\[ (2x^2 - 3y^2) -8x -42 +12y+36=0 \] replace the stuff in parens with \[ 2x^2 - 3y^2= -12x+12y-22 \] to get \[ -12x+12y-22-8x -42 +12y+36=0 \] simplify that mess and solve for x

OpenStudy (phi):

if you can get rid of quadratics, then you should do so...

OpenStudy (anonymous):

i didnt derive a quadratic..

OpenStudy (anonymous):

its -20x + 24y - 28

OpenStudy (phi):

you mean -20x + 24y - 28=0 divide by 4 -5x + 6y -7=0 x= (6y-7)/5 now use that value of x in our cross product: \[ 2(x+3)(x-7) = 3(y+2)(y-6) \] and we will get a quadratic in y solve for y... you get 2 values. find x using x= (6y-7)/5 find the slope. and then the line

OpenStudy (anonymous):

2 (6y-7/5) equals to?? 12y - 14/5? am i rght?

OpenStudy (phi):

\[ 2(x+3)(x-7) = 3(y+2)(y-6) \\ 2\left( \frac{(6y-7)}{5} + 3\right)\left( \frac{(6y-7)}{5} -7\right)= 3(y+2)(y-6) \]

OpenStudy (anonymous):

this is so algebra okay..

OpenStudy (anonymous):

ill answer it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

((12y+4/5) +6) ((6y-7)/5 -7)= 3y^2 - 12y - 12..

OpenStudy (phi):

( (12y-14)/5 +6) ((6y-7)/5 -7)= 3y^2 - 12y - 36

OpenStudy (anonymous):

(12y -14/5) (6y-7/5) = 72y^2 - 84y - 84y + 98 / 25??

OpenStudy (anonymous):

my answer is 72y^2 - 84y - 84y + 98/ 25 - 84y - 98/5 + 36y- 42/5 - 42.. am i right? o did i missed a thing? a sign?

OpenStudy (phi):

72y^2 - 84y/5 - 84y/5 + 98 / 25 or (12y -14/5) (6y-7/5) = 72y^2 -168y/5 + 98/25 I did it differently.. \[ 2\left( \frac{(6y-7)}{5} + 3\right)\left( \frac{(6y-7)}{5} -7\right)= 3(y+2)(y-6) \\ 2\left( \frac{(6y-7+15)}{5} \right)\left( \frac{6y-7-35}{5} \right)= 3(y+2)(y-6) \\ 2\left( \frac{6y+8}{5} \right)\left( \frac{6y-42}{5} \right)= 3(y+2)(y-6) \\ 2\cdot 2\left( \frac{3y+4}{5} \right)\cdot 6 \cdot \left( \frac{y-7}{5} \right)= 3(y+2)(y-6) \\ 24 \frac{(3y+4)(y-7)}{25}= 3(y^2 -4y -12) \\ 8 (3y+4)(y-7)= 75(y^2 -4y -12) \]

OpenStudy (phi):

the very last line in the above post should have 25 on the right side (not 75) I divided both sides by 3 \[ 8 (3y+4)(y-7)=25(y^2 -4y -12) \\ 8 (3y^2 -17y-28) =25y^2 - 100y -300 \\ 24y^2 - 136y -224= 25y^2 - 100y -300 \] now simplify

OpenStudy (anonymous):

wait. can i ask you about how did you derived 8?

OpenStudy (anonymous):

got it now. 24/3

OpenStudy (phi):

yes. and then multiply both sides by 25 ... and I forgot I did 3/3 on the right side

OpenStudy (anonymous):

i got 2 and - 38. FINALLY HOO. THANKS TO YOOOOOOU

OpenStudy (phi):

still some work to do... but the nasty stuff is over with...

OpenStudy (anonymous):

can i ask you something? about getting the dy/dx of xy=18?

OpenStudy (phi):

xy=18 you can solve for y \[ y = 18 x^{-1} \] \[ \frac{dy}{dx} = -18x^{-2} \]

OpenStudy (anonymous):

raised to -1. why??

OpenStudy (phi):

or use implicit differentiation x dy/dx + y dx/dx = 0 dy/dx = -y/x if you like, you can use y = 18/x and get dy/dx = -18/x^2

OpenStudy (phi):

raised to -1. why?? \[ \frac{1}{x}= x^{-1} \] right ? taking the derivative of a power is easy, so I change 1/x to x^(-1)

OpenStudy (anonymous):

squared?

OpenStudy (phi):

are you asking how to take the derivative of x^n ? \[ d x^n = n x^{n-1} dx \]

OpenStudy (anonymous):

okay!! got it now..

OpenStudy (anonymous):

thank you so much for the help, patience. yes and everything. need to finish my assignment. 2 more items to go. hope i can do this :) God bless you..